OpenStudy (kitkat16):

Help please! Alice works at a health care facility. She has measured the hemoglobin levels of 200 people. The data follows a normal distribution with a mean of 14 g/dL and a standard deviation of 1. How many people have hemoglobin levels less than 13? How many people have hemoglobin levels greater than 14?

OpenStudy (kitkat16):

so its 68% with a standard deviation of one

OpenStudy (kitkat16):

200-68% would it be 64 below and 64 greater? Help please?

OpenStudy (pinklion23):

you still need help with this?

OpenStudy (pinklion23):

http://inquus.org/study#/updates/54a22656e4b054f0c3b3ce21 this link might help you idk this one is confusing me lol :) hope it helped and if it did can i have a medal that way i know i was right?

OpenStudy (anonymous):

With 200 people in the sample, according to this distribution, 50% of the sample fall to either side of the mean. You can use this fact to answer the second question right away. The first question asks for the proportion of people with levels less than 13, which can be expressed as a probability: P(X<13)=P(X−141<13−141)=P(Z<−1).gives you the proportion of people with levels less than 13, so multiplying this proportion by 200 will tell you how many people in the sample have these levels. you may do the same for the second question

OpenStudy (kitkat16):

I still don't understand.

OpenStudy (kitkat16):

do I minus and add the 68% or is that wrong?

OpenStudy (amistre64):

since 14 is the mean, the mean of a normal distribution splits that area in half ...50% below and 50% above 13 is 1sd below the mean ..... your 68% idea is from the empirical rule. from 1 sd below the mean to 1 sd above the mean (13 to 15) 68% of the data fits in there 13 15 |---------|-----------------|----------| a b c d 68% so the total from a to d is 100% we want the value below 13, so from a to b (a to d) - (b to c) 100 - 68 = 32 32 is split between (a to b) and (c to d); divide it in half 13 15 |---------|-----------------|----------| a b c d 16% 68% 16%

OpenStudy (kitkat16):

thanks!!!